OpenStudy (anonymous):

Willie’s weight is 2 pounds less than twice his son’s weight. Willie weighs 174 pounds. Which of the following equations can be used to find the son’s weight? a. 174 = 2 + 2x c. 174 = 2x - 2 b. 174 = 2 - 2x d. 174 = 3x -2

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

a. 174 = 2 + 2x c. 174 = 2x - 2 b. 174 = 2 - 2x d. 174 = 3x -2

OpenStudy (anonymous):

there is no c

OpenStudy (anonymous):

yess there is

OpenStudy (anonymous):

c. 174 = 2x - 2

OpenStudy (anonymous):

oh wow hahaha my bad since theyr spacing was little i didnt notice